# Collision Avoidance and Geofencing for Fixed-wing Aircraft with Control Barrier Functions

### v2

@@ I Introduction
Safe behavior is of utmost importance for aerial vehicles due to the severity of consequences in case of an incidental failure. Thus, on-board control systems must satisfy strict safety constraints while operating aircraft during flight. Safety constraints may span a wide range of criteria, including bounds on the aircraft’s flight envelope and control surfaces (envelope protection), altitude, speed, acceleration, load factors, orientation, angular rates, angle of attack, control surface deflections and their rates, as well as bounds on an aircraft’s trajectory (safe navigation) that may include geofences (keep-in or keep-out zones) and collision avoidance with the ground or other air vehicles. Such strict safety specifications necessitate control systems that are designed in a safety-critical fashion, preferably with guarantees or certificates of safe behavior under certain operating conditions. Recently, this has led to the idea of run-time assurance (RTA) systems [1] that serve as an added module to the aircraft’s nominal flight controller and intervene whenever necessary to avoid the violation of safety. How to design control laws for use in RTA, however, is still an open problem.
@@
Safety has long been of interest in the aerospace control literature. Many studies have focused on safe aircraft navigation, where the overall motion is controlled in scenarios like collision avoidance, trajectory tracking and geofencing. Approaches for collision avoidance include the use of potential fields [2, 3], velocity obstacles [4, 5], control barrier functions (CBFs) [6], controlled invariant sets [7], and reinforcement learning [8, 9]. Notably, CBFs demonstrated success in collision avoidance on quadrotors in simulation [10] and experiments [11], as well as on fixed-wing aircraft in the context of probabilistic safety certificates [12], learning-based [13] and data-driven CBFs [14], and multi-aircraft control [15]. Trajectory tracking by fixed-wing aircraft also leveraged CBF theory to keep path following errors within prescribed bounds, via higher-order CBFs [16], barrier Lyapunov functions [17], and robust CBFs [18]. Moreover, geofencing—where the aircraft is navigated to avoid a restricted territory beyond a “geofence” like the airspace of a country or private property—was also addressed by CBFs with success on quadrotors [19, 20]. However, a general CBF-based RTA framework for the safe navigation of fixed-wing aircraft, including simultaneous collision avoidance and geofencing, is yet to be developed.
@@
In this paper, we propose a safety-critical RTA system for fixed-wing aircraft using CBFs, as summarized by Fig. 1. Specifically, we accomplish simultaneous collision avoidance and geofencing in 3D space on a nonlinear kinematic model of an aircraft. First, we encode multiple safety constraints (related to collision avoidance and multiple geofence boundaries) into a single CBF candidate. Second, we establish how to construct CBFs for cascaded aircraft dynamics. We introduce and compare three approaches: higher-order CBFs, backstepping-based CBFs, and model-free CBFs. Finally, we demonstrate the behavior of the proposed RTA by simulations of the kinematic model and a high-fidelity dynamical model. To the best of our knowledge, this is the first application of CBFs in the context of simultaneous collision avoidance and geofencing by fixed-wing aircraft with cascaded nonlinear dynamics.
@@
The paper is structured as follows. Section II describes the nonlinear model of fixed-wing aircraft kinematics. Section III provides background on CBFs. Section IV introduces the main contributions by establishing and simulating the proposed CBF-based RTA systems. Section V closes with conclusions.
@@ II Modeling the Aircraft’s Motion
First, we introduce a kinematic model, called 3D Dubins model, that describes the motion of fixed-wing aircraft. Kinematic models are popular in navigation and path planning [21, 22]. This model is derived in Appendix A from six-degrees-of-freedom dynamics with certain simplifying assumptions. The main assumption is that the angles of attack, $\alpha$, and sideslip, $\beta$, are negligible. In this paper, we use this model for safety-critical controller synthesis, to design acceleration and angular velocity commands for the aircraft in a provably safe fashion.
@@
$\begin{split}\dot{n}&=V_{\rm T}{\rm c}\psi{\rm c}\theta,\\ \dot{e}&=V_{\rm T}{\rm s}\psi{\rm c}\theta,\\ \dot{d}&=-V_{\rm T}{\rm s}\theta,\\ \dot{\phi}&=P+{\rm s}\phi{\rm t}\theta Q+{\rm c}\phi{\rm t}\theta R,\\ \dot{\theta}&={\rm c}\phi Q-{\rm s}\phi R,\\ \dot{\psi}&=\frac{{\rm s}\phi}{{\rm c}\theta}Q+\frac{{\rm c}\phi}{{\rm c}\theta}R,\\ \dot{V}_{\rm T}&=A_{\rm T},\end{split}$ (1)
where ${\rm s}$, ${\rm c}$, ${\rm t}$ stand for $\sin$, $\cos$, $\tan$ and:
$R=\frac{g_{\rm D}}{V_{\rm T}}{\rm s}\phi{\rm c}\theta.$ (2)
@@
$\mathbf{r}=\begin{bmatrix}n\\ e\\ d\end{bmatrix},\quad\boldsymbol{\zeta}=\begin{bmatrix}V_{\rm T}\\ \theta\\ \psi\end{bmatrix},\quad\mathbf{v}(\boldsymbol{\zeta})=\begin{bmatrix}V_{\rm T}{\rm c}\theta{\rm c}\psi\\ V_{\rm T}{\rm c}\theta{\rm s}\psi\\ -V_{\rm T}{\rm s}\theta\end{bmatrix}.$ (4)
With these, the dynamics (1)-(2) form a cascaded structure:
$\begin{split}\dot{\mathbf{r}}&=\mathbf{v}(\boldsymbol{\zeta}),\\ \dot{\boldsymbol{\zeta}}&=\mathbf{f}_{\boldsymbol{\zeta}}(\boldsymbol{\zeta},\phi,R,A_{\rm T},Q),\quad R=\rho(\boldsymbol{\zeta},\phi),\\ \dot{\phi}&=f_{\phi}(\boldsymbol{\zeta},\phi,R,Q,P),\end{split}$ (5)
that highlights the interdependence of states and inputs. Specifically, the evolution of the position $\mathbf{r}$ is given by the velocity-related states $\boldsymbol{\zeta}$. The evolution of $\boldsymbol{\zeta}$ depends on the remaining state $\phi$, the state-dependent turning rate $R$, and two control inputs $A_{\rm T}$ and $Q$. Finally, the evolution of the last state $\phi$ involves the third control input $P$. Overall, the dynamics have a 3-level cascaded structure, where the control inputs enter at the 2nd and 3rd level (with an auxiliary algebraic formula for $R$). Importantly, the right-hand side expressions:
$\mathbf{f}_{\boldsymbol{\zeta}}(\boldsymbol{\zeta},\phi,R,A_{\rm T},Q)$ $=\begin{bmatrix}A_{\rm T}\\ {\rm c}\phi Q-{\rm s}\phi R\\ \frac{{\rm s}\phi}{{\rm c}\theta}Q+\frac{{\rm c}\phi}{{\rm c}\theta}R\end{bmatrix},\quad\rho(\boldsymbol{\zeta},\phi)=\frac{g_{\rm D}}{V_{\rm T}}{\rm s}\phi{\rm c}\theta,$ (9)
$f_{\phi}(\boldsymbol{\zeta},\phi,R,Q,P)$ $=P+{\rm s}\phi{\rm t}\theta Q+{\rm c}\phi{\rm t}\theta R$ (10)
@@
$\begin{split}\mathbf{f}(\mathbf{x})&=\begin{bmatrix}\mathbf{v}(\boldsymbol{\zeta})\\ \mathbf{f}_{\boldsymbol{\xi}}(\mathbf{x})\\ 0\end{bmatrix},\quad\mathbf{g}(\mathbf{x})=\begin{bmatrix}\hfil\mathbf{0}_{3\times 3}\hfil\\ \hfil\mathbf{g}_{\boldsymbol{\xi}}(\mathbf{x})\hfil\\ 1&0&0\end{bmatrix},\end{split}$ (12)
@@
$\begin{split}\mathbf{f}_{\boldsymbol{\xi}}(\mathbf{x})=\frac{g_{\rm D}}{V_{\rm T}}\begin{bmatrix}{\rm s}\phi{\rm c}\phi{\rm s}\theta\\ -({\rm s}\phi)^{2}{\rm c}\theta\\ {\rm s}\phi{\rm c}\phi\end{bmatrix},\quad\mathbf{g}_{\boldsymbol{\xi}}(\mathbf{x})=\begin{bmatrix}0&1&{\rm s}\phi{\rm t}\theta\\ 0&0&{\rm c}\phi\\ 0&0&\frac{{\rm s}\phi}{{\rm c}\theta}\end{bmatrix}.\end{split}$ (13)
In what follows, we synthesize controllers for the aircraft model (1)-(2), with the end goal of formal safety guarantees w.r.t. collision avoidance and geofencing. In particular, we design safe controllers for the control-affine form (11) of the model by using control barrier functions constructed based on the cascaded structure (5) of the dynamics.
@@ III-B Safety Filters
Based on Theorem 1, controllers can be designed for (14) in a safety-critical fashion by enforcing (19) during controller synthesis. Such safety-critical control designs are often established by so-called safety filters that modify a desired but not necessarily safe controller ${\mathbf{k}_{\rm d}:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{m}}$ to a safe controller $\mathbf{k}$. In the context of aircraft control, the desired controller $\mathbf{k}_{\rm d}$ may be a flight controller that tracks a desired trajectory, such as the controller in Appendix B. This can be modified for safety, for example, by the optimization problem:
@@
Importantly, the optimization problem (20) can be solved in closed form. By factorizing $\boldsymbol{\Gamma}$ as ${\boldsymbol{\Gamma}=\mathbf{W}^{-\top}\mathbf{W}^{-1}}$ with positive definite ${\mathbf{W}\in\mathbb{R}^{m\times m}}$, (20) transforms into:
@@
Then (21) can be solved as follows, providing an explicit, easy-to-evaluate safety filter [25]:
@@ III-C Compositions of Safety Constraints
Combinations of set unions and intersections can be addressed by the recursive applications of formulas (27) and (31). Overall, these formulas provide a systematic way of encoding complex compositions of safety constraints into a single CBF candidate. This will be leveraged below in the context of simultaneous collision avoidance and geofencing where the geofence is composed of multiple boundaries.

### v3

@@ I Introduction
Safe behavior is of utmost importance for aerial vehicles due to the severity of consequences in case of an incidental failure. Thus, on-board control systems must satisfy strict safety constraints while operating aircraft during flight. Safety constraints may span a wide range of criteria, including bounds on the aircraft’s flight envelope and control surfaces (envelope protection), altitude, speed, acceleration, load factors, orientation, angular rates, angle of attack, control surface deflections and their rates, as well as bounds on an aircraft’s trajectory (safe navigation) that may include geofences (keep-in or keep-out zones) and collision avoidance with the ground or other air vehicles. Such strict safety specifications necessitate control systems that are designed in a safety-critical fashion, preferably with guarantees or certificates of safe behavior under certain operating conditions. Recently, this has led to the idea of run-time assurance (RTA) systems [1] that serve as an added module to the aircraft’s nominal flight controller and intervene whenever necessary to avoid the violation of safety. RTA is often used to supervise complex control systems whose safe behavior is difficult to guarantee or verify, such as learning-based controllers [2, 3]. RTA systems showed promising results in indoor flight with quadrotors [4], navigation of unmanned fixed-wing aircraft along safe corridors [5], neural network-based aircraft taxiing application [6], and aerial refueling task in naval aviation [7]. How to design safe control laws for use in RTA, however, is still an open problem.
@@
Safety has long been of interest in the aerospace control literature. Many studies have focused on safe aircraft navigation, where the overall motion is controlled in scenarios like collision avoidance, trajectory tracking, and geofencing. Approaches include the use of potential fields [8, 9], velocity obstacles [10], control barrier functions (CBFs) [11], controlled invariant sets [12], and reinforcement learning [13, 14]. Notably, CBFs provide provable safety guarantees as opposed to most learning-based methods. Furthermore, some CBF formulations can be used to generate control invariant sets [15], while CBFs can even be constructed from artificial potential fields [16] and velocity obstacles [17]. Most CBF-based controllers respond to the current state of the system without a preview of future trajectories. This may lead to suboptimal behavior compared to other methods with preview, such as motion planning algorithms or model predictive control, but it also makes CBF-based controllers easy to compute.
@@
CBFs demonstrated success in collision avoidance on quadrotors in simulation [18, 19] and experiments [16], as well as on fixed-wing aircraft in the context of probabilistic safety certificates [20], learning-based [21] and data-driven CBFs [22], and multi-aircraft control [23]. Trajectory tracking by fixed-wing aircraft also leveraged CBF theory to keep path following errors within prescribed bounds, via high-order CBFs [24], barrier Lyapunov functions [25], and robust CBFs [26]. Moreover, geofencing—where the aircraft is navigated to avoid a restricted territory beyond a “geofence” like the airspace of a country or private property—was also addressed by CBFs with success on quadrotors [27, 28]. A CBF-based fixed-wing aircraft RTA system for simultaneous collision avoidance and geofencing is yet to be developed.
Notably, [29] addressed simultaneous collision avoidance and geofencing without CBFs, for tactical unmanned aircraft considering multiple intruder aircraft, fixed obstacles, no-fly zones, and bad weather areas. An optimal control problem was formulated with risk assessment and probabilistic safety constraints. In our present work, we instead focus on deterministic safety constraints, and develop a CBF-based RTA system for simultaneous collision avoidance and geofencing. We use a reduced-order model of the aircraft dynamics to formulate the CBF and synthesize a controller in closed form that provides formal guarantees of safety. The approach of developing safety-critical controllers via reduced-order models and CBFs has been successful on a variety of autonomous systems, including legged, wheeled, and flying robots, manipulators, and heavy-duty trucks [30]. Now we seek to extend this approach to establish RTA for fixed-wing aircraft.
@@
In this paper, we propose a safety-critical RTA system for fixed-wing aircraft using CBFs, as summarized by Fig. 1. Specifically, we accomplish simultaneous collision avoidance and geofencing in 3D space on a nonlinear kinematic model of an aircraft. To the best of our knowledge, this is the first application of CBFs to simultaneous collision avoidance and geofencing for fixed-wing aircraft. First, we encode multiple safety constraints (related to collision avoidance and multiple geofence boundaries) into a single CBF candidate. Second, we establish how to construct CBFs for nonlinear aircraft dynamics. We introduce and compare three approaches: high-order CBFs, backstepping-based CBFs, and model-free CBFs. We highlight that a careful CBF construction is required to leverage all control inputs and make the aircraft both accelerate, pitch, and roll for safe collision avoidance and geofencing. This challenge can be addressed, for example, via backstepping, and this work is the first application of CBF backstepping to fixed-wing aircraft. Finally, we demonstrate the behavior of the proposed RTA by simulations of the kinematic model and also a high-fidelity dynamical model.
@@
The paper is structured as follows. Section II describes the nonlinear model of fixed-wing aircraft kinematics. Section III provides background on CBFs. Section IV introduces the main contributions by establishing and simulating the proposed CBF-based RTA system. Section V closes with conclusions.
@@ II Modeling the Aircraft’s Motion
First, we introduce a kinematic model, called 3D Dubins model, that describes the motion of fixed-wing aircraft. Kinematic models are popular in navigation and path planning [31, 32]. This model is derived in Appendix A from six-degrees-of-freedom dynamics using certain assumptions (cf. Assumption 3). The main assumption is that the angles of attack and sideslip are negligible. We use this model for safety-critical controller synthesis, to design acceleration and angular velocity commands for the aircraft in a provably safe fashion.
@@
$\begin{split}\dot{n}&=V_{\rm T}{\rm c}_{\psi}{\rm c}_{\theta},\\ \dot{e}&=V_{\rm T}{\rm s}_{\psi}{\rm c}_{\theta},\\ \dot{d}&=-V_{\rm T}{\rm s}_{\theta},\\ \dot{\phi}&=P+{\rm s}_{\phi}{\rm t}_{\theta}Q+{\rm c}_{\phi}{\rm t}_{\theta}R,\\ \dot{\theta}&={\rm c}_{\phi}Q-{\rm s}_{\phi}R,\\ \dot{\psi}&=\frac{{\rm s}_{\phi}}{{\rm c}_{\theta}}Q+\frac{{\rm c}_{\phi}}{{\rm c}_{\theta}}R,\\ \dot{V}_{\rm T}&=A_{\rm T},\end{split}$ (1)
where ${\rm s}_{(.)}$, ${\rm c}_{(.)}$, ${\rm t}_{(.)}$ stand for $\sin(.)$, $\cos(.)$, $\tan(.)$ and:
$R=\frac{g_{\rm D}}{V_{\rm T}}{\rm s}_{\phi}{\rm c}_{\theta},$ (2)
with $g_{\rm D}$ being the gravitational acceleration.
@@
$\mathbf{r}=\begin{bmatrix}n\\ e\\ d\end{bmatrix},\quad\boldsymbol{\zeta}=\begin{bmatrix}V_{\rm T}\\ \theta\\ \psi\end{bmatrix},\quad\mathbf{v}(\boldsymbol{\zeta})=\begin{bmatrix}V_{\rm T}{\rm c}_{\theta}{\rm c}_{\psi}\\ V_{\rm T}{\rm c}_{\theta}{\rm s}_{\psi}\\ -V_{\rm T}{\rm s}_{\theta}\end{bmatrix}.$ (4)
With these, the dynamics (1)-(2) become:
$\begin{split}\dot{\mathbf{r}}&=\mathbf{v}(\boldsymbol{\zeta}),\\ \dot{\boldsymbol{\zeta}}&=\mathbf{f}_{\boldsymbol{\zeta}}(\boldsymbol{\zeta},\phi,R,A_{\rm T},Q),\quad R=\rho(\boldsymbol{\zeta},\phi),\\ \dot{\phi}&=f_{\phi}(\boldsymbol{\zeta},\phi,R,Q,P).\end{split}$ (5)
These dynamics form a cascaded structure, where the evolution of higher-level states depends on lower-level states. Specifically, the evolution of the position $\mathbf{r}$ is given by the velocity-related states $\boldsymbol{\zeta}$. The evolution of $\boldsymbol{\zeta}$ depends on the remaining state $\phi$, the state-dependent turning rate $R$, and two control inputs $A_{\rm T}$ and $Q$. Finally, the evolution of the last state $\phi$ involves the third control input $P$. Overall, the dynamics have a 3-level cascaded structure, where the control inputs enter at the 2nd and 3rd level (with an auxiliary algebraic formula for $R$). Importantly, the right-hand side expressions:
$\mathbf{f}_{\boldsymbol{\zeta}}(\boldsymbol{\zeta},\phi,R,A_{\rm T},Q)$ $=\begin{bmatrix}A_{\rm T}\\ {\rm c}_{\phi}Q-{\rm s}_{\phi}R\\ \frac{{\rm s}_{\phi}}{{\rm c}_{\theta}}Q+\frac{{\rm c}_{\phi}}{{\rm c}_{\theta}}R\end{bmatrix},\quad\rho(\boldsymbol{\zeta},\phi)=\frac{g_{\rm D}}{V_{\rm T}}{\rm s}_{\phi}{\rm c}_{\theta},$
$f_{\phi}(\boldsymbol{\zeta},\phi,R,Q,P)$ $=P+{\rm s}_{\phi}{\rm t}_{\theta}Q+{\rm c}_{\phi}{\rm t}_{\theta}R$ (6)
@@
[…]
@@
$\begin{split}\mathbf{f}_{\boldsymbol{\xi}}(\mathbf{x})=\frac{g_{\rm D}}{V_{\rm T}}\begin{bmatrix}{\rm s}_{\phi}{\rm c}_{\phi}{\rm s}_{\theta}\\ -{\rm s}_{\phi}^{2}{\rm c}_{\theta}\\ {\rm s}_{\phi}{\rm c}_{\phi}\end{bmatrix},\quad\mathbf{g}_{\boldsymbol{\xi}}(\mathbf{x})=\begin{bmatrix}0&1&{\rm s}_{\phi}{\rm t}_{\theta}\\ 0&0&{\rm c}_{\phi}\\ 0&0&\frac{{\rm s}_{\phi}}{{\rm c}_{\theta}}\end{bmatrix}.\end{split}$ (9)
In what follows, we synthesize controllers for the aircraft model (1)-(2), with the end goal of formal safety guarantees w.r.t. collision avoidance and geofencing. In particular, we design safe controllers for the control-affine form (7) of the model by using control barrier functions constructed based on the cascaded structure (5) of the dynamics. When designing the controller, we consider that the position $\mathbf{r}$ of the aircraft is safe at time $t$ if the pair $(\mathbf{r},t)$ is inside a safe set ${\mathcal{S}\subset\mathbb{R}^{3}\times\mathbb{R}_{\geq 0}}$ defined by the collision avoidance and geofencing constraints (see details in Section IV).
Problem Statement.
Consider system (7) and design a controller ${\mathbf{k}:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{m}}$, ${\mathbf{u}=\mathbf{k}(\mathbf{x},t)}$ that guarantees safety w.r.t. collision avoidance and geofencing. Specifically, the controller must ensure that the trajectory $\mathbf{r}(t)$ of the closed-loop system is maintained within a prescribed set $\mathcal{S}$: ${(\mathbf{r}(t),t)\in\mathcal{S}}$, ${\forall t\geq 0}$ for initial positions satisfying ${(\mathbf{r}_{0},0)\in\mathcal{S}}$.
@@ III-A Theoretical Background
Note that if the derivative of $h$ is independent of the control input $\mathbf{u}$, i.e., ${\frac{\partial h}{\partial\mathbf{x}}(\mathbf{x},t)\mathbf{g}(\mathbf{x})\equiv\mathbf{0}}$ in (13), then $h$ cannot be used directly to synthesize safe controllers. In such cases, methods like high-order CBFs [35, 36] or backstepping [37] can be applied to construct a CBF from $h$ and use it for control. These methods are discussed further in Section IV. The main idea behind high-order CBFs is introducing ${\dot{h}+\alpha(h)}$, that occurs in (15), as a high-order (or extended) CBF:
$h_{\rm e}(\mathbf{x},t)=\frac{\partial h}{\partial t}(\mathbf{x},t)+\frac{\partial h}{\partial\mathbf{x}}(\mathbf{x},t)\mathbf{f}(\mathbf{x})+\alpha\big{(}h(\mathbf{x},t)\big{)},$ (16)
whose 0-superlevel set is $\mathcal{C}_{\rm e}$, cf. (12). If the extended CBF $h_{\rm e}$ is maintained nonnegative, then safety can be guaranteed based on (15), similar to the case when $h$ is a CBF.
Lemma 2 ([36]).
If $h_{\rm e}$ is a CBF for (10) on $\mathcal{C}_{\rm e}$, then any locally Lipschitz continuous controller $\mathbf{k}$ that satisfies:
$\dot{h}_{\rm e}\big{(}\mathbf{x},t,\mathbf{k}(\mathbf{x},t)\big{)}\geq-\alpha\big{(}h_{\rm e}(\mathbf{x},t)\big{)}$ (17)
for all ${(\mathbf{x},t)\in\mathcal{C}\cap\mathcal{C}_{\rm e}}$ renders (11) safe w.r.t. $\mathcal{C}\cap\mathcal{C}_{\rm e}$.
@@ III-B Safety Filters
Based on Lemma 1, controllers can be designed for (10) in a safety-critical fashion by enforcing (15) during controller synthesis. Such safety-critical control designs are often established by so-called safety filters that modify a desired but not necessarily safe controller ${\mathbf{k}_{\rm d}:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{m}}$ to a safe controller $\mathbf{k}$. In the context of aircraft control, the desired controller $\mathbf{k}_{\rm d}$ may be a flight controller that tracks a desired trajectory, such as the controller in Appendix B. This can be modified for safety, for example, by the optimization problem:
@@
We remark that the formulation above does not include input constraints (as ${\mathbf{u}\in\mathbb{R}^{m}}$). To enforce input limits, one could saturate the controller, or add input constraints to (18) and relax the safety constraint to maintain feasibility. However, these methods lose formal safety guarantees. Formally addressing input constraints while maintaining safety requires more advanced CBF approaches, such as those in [15, 38, 39, 40, 41]. These methods are beyond the scope of this paper and will be explored as future work. Instead, the controllers proposed here are tuned so that the resulting control inputs remain within reasonable bounds in simulations.
Importantly, the optimization problem (18) is feasible if $h$ is indeed a CBF that satisfies (14), and it can be solved in closed form. By factorizing $\boldsymbol{\Gamma}$ as ${\boldsymbol{\Gamma}=\mathbf{W}^{-\top}\mathbf{W}^{-1}}$ with positive definite ${\mathbf{W}\in\mathbb{R}^{m\times m}}$, (18) transforms into:
@@
Then (19) can be solved as follows [42]:
@@
Since the safety filter (21) is given explicitly, it is fast and easy to evaluate it, which is an advantage of using CBFs.
@@ III-C Compositions of Safety Constraints
Formula (25) includes the evaluation of $N_{\rm c}$ exponential functions, hence its computation time increases with the number $N_{\rm c}$ of constraints. Yet, a significant benefit of constructing a single CBF by (25) is that it enables the use of closed-form safety filters like (21) that are easy to compute. Without such composition, safety filters would require solving optimization problems like (18) with $N_{\rm c}$ number of constraints, which would ultimately lead to more computation especially if $N_{\rm c}$ is large. Further details on the properties of the CBF composition method in (25) can be found in [43].
@@
Combinations of set unions and intersections can be addressed by the recursive applications of formulas (25) and (29). Note that (14) must hold for $h$ to be a valid CBF. This requires a control-sharing property [45, 46] that the individual barriers $h_{i}$ are compatible and do not lead to contradicting constraints for the control input. Establishing the conditions for obtaining a valid CBF after a general composition is still an open problem, and initial results can be found in [43]. Overall, formulas (25) and (29) provide a systematic way of encoding complex compositions of safety constraints into a single CBF candidate. This will be leveraged below in the context of simultaneous collision avoidance and geofencing where the geofence is composed of multiple boundaries.